# A Bias-Free Training Paradigm for More General AI-generated Image Detection

### v1

@@ Abstract
Successful forensic detectors can produce excellent results in supervised learning benchmarks but struggle to transfer to real-world applications. We believe this limitation is largely due to inadequate training data quality. While most research focuses on developing new algorithms, less attention is given to training data selection, despite evidence that performance can be strongly impacted by spurious correlations such as content, format, or resolution. A well-designed forensic detector should detect generator specific artifacts rather than reflect data biases. To this end, we propose B-Free, a bias-free training paradigm, where fake images are generated from real ones using the conditioning procedure of stable diffusion models. This ensures semantic alignment between real and fake images, allowing any differences to stem solely from the subtle artifacts introduced by AI generation. Through content-based augmentation, we show significant improvements in both generalization and robustness over state-of-the-art detectors and more calibrated results across 27 different generative models, including recent releases, like FLUX and Stable Diffusion 3.5. Our findings emphasize the importance of a careful dataset curation, highlighting the need for further research in dataset design. Code and data will be publicly available at https://grip-unina.github.io/B-Free/.
@@ 1 Introduction
The rise of generative AI has revolutionized the creation of synthetic content, enabling easy creation of high-quality sophisticated context, even for individuals without deep technical expertise. Thanks to user-friendly interfaces and pretrained models, users can create high-quality synthetic content such as text, images, music, and videos through simple inputs or prompts [50]. This accessibility has democratized content creation, enabling professionals in fields like design, marketing, and entertainment to leverage AI for creative purposes. However, this raises concerns about potential misuse, such as the creation of deepfakes, misinformation, and challenges related to intellectual property and content authenticity [16, 4, 25].
@@
Figure 2: Forensic detectors can exhibit opposite behaviors depending on their training dataset. The four plots show the prediction distributions for three CLIP-based detectors, UnivFD [31], FatFormer [27] and RINE [21], and the proposed one. The fake images (SD-XL or DALL-E 3) are generated from images of a single dataset (RAISE on top, COCO on the bottom) and tested only against real images of the same dataset (Synthbuster [2] and the test dataset from [11]). We observe that for the same detector (e.g., RINE) and the same fake-image generator (e.g., DALL-E 3) the score distributions can vary significantly depending on the dataset used, going from real (left of the dotted line) to fake (right of the dotted line) or vice versa. This is likely due to the presence of biases in the training set that heavily impact the detector prediction. Our detector, on the other hand, shows consistent and correct results in all cases.
@@
Key challenges for current GenAI image detectors include generalization — detecting synthetic generators not present in the training set — and ensuring robustness against image impairments caused by online sharing, such as compression, resizing, and cropping [41]. In this context, large pre-trained vision-language models like CLIP [34] have demonstrated impressive resilience to these distribution shifts [31]. The success of these models in forensic applications suggests that pre-training on large and diverse datasets may be a promising path forward. An important aspect often overlooked in the current literature is selection of good task specific datasets to train or fine-tune such models, that primarily rely on hidden, unknown signatures of generative models [30, 47]. Indeed, it is important to guarantee that the detector decisions are truly based on generation-specific artifacts and not on possible dataset biases [42, 28, 7]. In fact, datasets used during the training and testing phases of forensic classifiers could be affected by different types of polarization.
@@
Format issues have been the Achilles’ heel of forensic detectors since at least 2013, when [5] recognized that a dataset for image tampering detection [13] included forged and pristine images compressed with different JPEG quality factors. Therefore, a classifier trained to discrminate tampered and pristine images may instead learn their different processing histories. This issue has been highlighted in [18] with reference to datasets of synthetic and real images. In fact, the former are often created in a lossless format (PNG), while the latter are typically compressed in lossy formats like JPEG. Again a classifier could learn coding inconsistencies instead of forensic clues. Likewise it could learn resampling artifacts, as it was recently shown in [35] - in this case a bias was introduced by resizing all the real images from the LAION dataset to same resolution, while keeping the fake ones unaltered.
@@
Forensic clues are subtle and often imperceptible to the human eye, making it easy to introduce biases when constructing the training and test sets, as well as the evaluation protocol. Semantic content itself can also represent a source of bias. For this reason, several recent proposals [2, 11, 3] take great care to include pairs of real and fake images characterized by the same prompts when building a training or test dataset. To gain better insights about the above issues, in Fig. 2 we show the performance of three SoTA CLIP-based approaches [31, 21, 27] in distinguishing real images from fake images generated by SD-XL and DALL-E 3. For each method we consider two settings: in the first case, real images come from the RAISE dataset [12] and fakes are generated starting from images of the same dataset. The second case uses COCO as source of reals instead of RAISE. FakeInversion [7] proposes an effective approach towards semantic alignment of training data using reverse image search to find matching reals, however fails to capture real image distribution after 2021.
@@
To mitigate potential dataset biases, in this work we propose a new training paradigm, B-Free, where we generate synthetic images using self-conditioned reconstructions of real images and incorporate augmented, inpainted variations. This approach helps prevent semantic bias and potential misalignment in coding formats. The model analyzes the whole image during training to avoid unnecessary cropping or resizing operations, the latter being particularly harmful since it could wash out the subtle low-level forensic clues [17]. Overall, we make the following contributions:
@@
  • We propose a large curated training dataset of 51k real and 309k fake images. Real images are sourced from COCO, while synthetic images are self-conditioned reconstructions using Stable Diffusion 2.1. This helps the detector to focus on artifacts related to the synthetic generation process avoiding content and coding related biases.
@@
  • We show that including proper content-based augmentation provides better calibrated results. This ensures that in-lab performance is closer to expected performance on real-world images shared over social networks.
@@
  • We study the effect of different distribution shifts and show that by leveraging a pre-trained large model trained on our dataset, it is possible to obtain a SoTA accuracy (87%) even on unseen new generators.
@@
Figure 3: Overview of existing ($a$, $b$, $c$) and proposed ($d$) strategies for building an aligned training dataset. Some methods try to match synthetic images to the corresponding real images by using class-based generation ($a$) or text-to-image generation with real images’ descriptions ($b$). In ($c$) real images are fed to an autoencoder to generate a reconstructed fake with the same content. Unlike in (c), in our approach ($d$) a self-conditioned fake is generated using diffusion steps, and we also add a content augmentation step.
@@ 2 Related Work
Other methods rely on datasets comprising images from a single diffusion-based generator, such as Latent Diffusion [10, 11, 7], Guided Diffusion [44] or Stable Diffusion [37, 22]. Prior work [11, 7] highlights the importance of aligning training and eval data in terms of semantic content. This choice allowed to better exploit the potential of fixed-pretraining CLIP features by strongly reducing the number of images needed for fine-tuning [11]. This strategy has the key merit of reducing the dataset content bias, thus allowing for better quality training, and is also adopted in other approaches both during training [1, 3] and test time to carry out a fairer evaluation [2].
@@
Table 1: Publicly available datasets we used for our evaluation.
@@ 3.1 Datasets
Training settings Synthbuster New Generators WildRF AVG
@@
AUC / bAcc (%) (b) paired by text -- 97.5 / 55.9 99.1 / 65.2 74.5 / 50.1 99.4 / 62.7 71.3 / 50.1 98.3 / 55.4 92.6 / 53.0 64.7 / 63.7 56.5 / 57.7 83.8 / 71.2 83.8 / 58.5
(c) reconstructed -- 100. / 99.4 100. / 99.1 77.7 / 52.7 97.0 / 72.3 99.3 / 89.3 94.5 / 67.0 98.4 / 84.3 98.1 / 94.7 75.5 / 75.3 93.3 / 88.4 93.4 / 82.3
(d) self-conditioned -- 99.8 / 96.9 99.7 / 96.6 89.4 / 58.1 97.9 / 82.5 100. / 99.1 89.5 / 57.6 98.9 / 89.6 95.6 / 90.6 79.3 / 75.5 94.2 / 89.0 94.4 / 83.6
@@
inpainted 100. / 96.8 100. / 96.8 81.2 / 65.0 99.1 / 95.1 100. / 96.8 94.3 / 84.9 99.6 / 95.9 99.0 / 87.5 81.0 / 70.5 96.7 / 86.1 95.1 / 87.5
inpainted+ 99.8 / 96.4 99.7 / 96.4 92.9 / 83.9 99.0 / 95.2 99.9 / 96.9 94.8 / 87.7 93.4 / 86.8 95.7 / 90.3 82.3 / 75.5 96.0 / 90.3 95.4 / 89.9
inpainted++ 99.9 / 96.8 99.8 / 96.6 83.1 / 71.0 97.4 / 92.2 99.5 / 96.7 92.9 / 85.0 94.7 / 89.5 97.7 / 92.5 83.2 / 78.8 95.8 / 89.5 94.4 / 88.9
@@
NLL / ECE (b) paired by text -- 2.21 / .427 1.41 / .342 4.83 / .496 1.45 / .364 5.04 / .498 2.10 / .429 3.06 / .460 1.83 / .300 2.66 / .366 1.41 / .253 2.60 / .393
(c) reconstructed -- 0.02 / .006 0.02 / .006 4.27 / .463 1.13 / .263 0.29 / .104 1.66 / .316 0.60 / .150 0.23 / .046 1.82 / .221 0.62 / .096 1.07 / .167
(d) self-conditioned -- 0.10 / .026 0.11 / .025 2.57 / .401 0.67 / .165 0.02 / .005 2.57 / .404 0.36 / .094 0.39 / .076 1.36 / .195 0.52 / .099 0.87 / .149
@@
inpainted 0.08 / .046 0.09 / .045 0.99 / .275 0.13 / .022 0.08 / .046 0.37 / .094 0.11 / .032 0.33 / .157 0.72 / .206 0.32 / .127 0.32 / .105
inpainted+ 0.09 / .039 0.10 / .036 0.44 / .099 0.13 / .023 0.08 / .042 0.34 / .065 0.40 / .077 0.28 / .067 0.70 / .155 0.27 / .054 0.28 / .066
inpainted++ 0.11 / .060 0.11 / .061 0.68 / .193 0.22 / .040 0.13 / .064 0.36 / .076 0.29 / .040 0.20 / .031 0.54 / .111 0.26 / .071 0.29 / .075
Table 2: Ablation study. We compare several forms of content alignment and content augmentation. Performance are in terms of AUC/Accuracy (top) and ECE/NLL (bottom). Note that all variants share a standard augmentation (blurring + JPEG compression) as proposed in [43]. For content alignment we consider the image pairing strategies described in Fig. 3: b) text-driven generation, c) reconstruction through autoencoder, and our proposal d) using self-conditioned images (through autoencoder and diffusion). For the last solution we test several forms of augmentation: a standard cutmix/mixup, and three proposed strategies based on inpainting: inpainted (same category and corresponding local version), inpainted+ (both same and different categories and corresponding local versions), and inpainted++ (like inpainted+ with further augmentation operations, such as scaling, cut-out, noise addition, and jittering).
3.2 Metrics
Most work on GenAI image detection measure performance by means of threshold-independent metrics, such as Area Under the Curve (AUC) or average precision (AP). These metrics indicate ideal classification performance, however the optimal separating threshold is not known and, quite often, the balanced accuracy at a fixed threshold (e.g. $0.5$) remains low, especially when there are significant differences between training and testing distributions [41]. Some papers address this problem by adjusting the threshold through a calibration procedure, assuming access to a few images from the synthetic generator under evaluation [43, 10, 31]. In a realistic situation, availability of such calibration images is uncertain.
In this work, to provide a comprehensive assessment of performance, we use both AUC and Accuracy at $0.5$, in addition we compute the Expected Calibration Error (ECE) and the Negative Log-Likelihood (NLL). ECE measures the ability of a model to provide prediction probabilities well aligned with the true probabilities. More precisely, we use the Binary ECE, which is the weighted average of the differences between the actual probability and the predicted probability across different bins [32]. Then, we use the balanced NLL [33], which evaluates the similarity between the distribution of the model’s predictions and the actual data distribution, penalizing both low confidence in the correct class and overconfidence in incorrect ones. More details on these metrics can be found in the supplementary material.

### v2

@@ Abstract
Successful forensic detectors can produce excellent results in supervised learning benchmarks but struggle to transfer to real-world applications. We believe this limitation is largely due to inadequate training data quality. While most research focuses on developing new algorithms, less attention is given to training data selection, despite evidence that performance can be strongly impacted by spurious correlations such as content, format, or resolution. A well-designed forensic detector should detect generator specific artifacts rather than reflect data biases. To this end, we propose B-Free, a bias-free training paradigm, where fake images are generated from real ones using the conditioning procedure of stable diffusion models. This ensures semantic alignment between real and fake images, allowing any differences to stem solely from the subtle artifacts introduced by AI generation. Through content-based augmentation, we show significant improvements in both generalization and robustness over state-of-the-art detectors and more calibrated results across 27 different generative models, including recent releases, like FLUX and Stable Diffusion 3.5. Our findings emphasize the importance of a careful dataset design, highlighting the need for further research on this topic. Code and data are publicly available at https://grip-unina.github.io/B-Free/.
@@ 1 Introduction
The rise of generative AI has revolutionized the creation of synthetic content, enabling easy production of high-quality sophisticated media, even for individuals without deep technical expertise. Thanks to user-friendly interfaces and pretrained models, users can create synthetic content such as text, images, music, and videos through simple inputs or prompts [54]. This accessibility has democratized content creation, enabling professionals in fields like design, marketing, and entertainment to leverage AI for creative purposes. However, this raises concerns about potential misuse, such as the creation of deepfakes, misinformation, and challenges related to intellectual property and content authenticity [17, 4, 27].
@@
Figure 2: Forensic detectors can exhibit opposite behaviors depending on their training dataset. The four plots show the prediction distributions for three ViT-based detectors, UnivFD [33], FatFormer [29] and RINE [22], and the proposed one. The fake images (SD-XL or DALL-E 3) are generated from images of a single dataset (RAISE on top, COCO on the bottom) and tested only against real images of the same dataset (Synthbuster [2] and the test dataset from [11]). We observe that for the same detector (e.g., RINE) and the same fake-image generator (e.g., DALL-E 3) the score distributions can vary significantly depending on the dataset used, going from real (left of the dotted line) to fake (right of the dotted line) or vice versa. This is likely due to the presence of biases in the training set that heavily impact the detector prediction. Our detector, on the other hand, shows consistent and correct results.
@@
Key challenges for current GenAI image detectors include generalization — detecting synthetic generators not present in the training set — and ensuring robustness against image impairments caused by online sharing, such as compression, resizing, and cropping [44]. In this context, large pre-trained vision-language models like CLIP [37] have demonstrated impressive resilience to these distribution shifts [33]. The success of these models in forensic applications suggests that pre-training on large and diverse datasets may be a promising path forward. An important aspect often overlooked in the current literature is the selection of good datasets to train or fine-tune such models that primarily rely on hidden, unknown signatures of generative models [32, 50]. Indeed, it is important to guarantee that the detector decisions are truly based on generation-specific artifacts and not on possible dataset biases [45, 30, 7]. In fact, datasets used during the training and testing phases of forensic classifiers could be affected by different types of polarization.
@@
Format issues have been the Achilles’ heel of forensic detectors since at least 2013, when [5] recognized that a dataset for image tampering detection [14] included forged and pristine images compressed with different JPEG quality factors. Therefore, a classifier trained to discrminate tampered and pristine images may instead learn their different processing histories. This issue has been highlighted in [19] with reference to datasets of synthetic and real images. In fact, the former are often created in a lossless format (PNG), while the latter are typically compressed in lossy formats like JPEG. Again a classifier could learn coding inconsistencies instead of forensic clues. Likewise it could learn resampling artifacts, as it was recently shown in [38] - in this case a bias was introduced by resizing all the real images from the LAION dataset to the same resolution, while keeping the fake ones unaltered.
@@
Forensic clues are subtle and often imperceptible to the human eye, making it easy to introduce biases when constructing the training and test sets, as well as the evaluation protocol. Semantic content itself can also represent a source of bias. For this reason, several recent proposals [2, 11, 3] take great care to include pairs of real and fake images characterized by the same prompts when building a training or test dataset. To gain better insights about the above issues, in Fig. 2 we show the performance of three SoTA ViT-based approaches [33, 22, 29] in distinguishing real images from fake images generated by SD-XL and DALL-E 3. For each method we consider two settings: in the first case, real images come from the RAISE dataset [12] and fakes are generated starting from images of the same dataset. The second case uses COCO as source of reals instead of RAISE. We note an inconsistent behavior of SOTA forensic detectors on the same synthetic generator which can be caused by the presence of biases during training. FakeInversion [7] proposes an effective approach towards semantic alignment of training data using reverse image search to find matching reals, but fails to capture real image distribution after 2021.
@@
To mitigate potential dataset biases, in this work we propose a new training paradigm, B-Free, where synthetic images are generated using self-conditioned reconstructions of real images and incorporate augmented, inpainted variations. This approach helps to prevent semantic bias and potential misalignment in coding formats. Furthermore, the model avoids resizing operations that can be particularly harmful by washing out the subtle low-level forensic clues [18]. Overall, we make the following contributions:
@@
  • We propose a large curated training dataset of 51k real and 309k fake images. Real images are sourced from COCO, while synthetic images are self-conditioned generations using Stable Diffusion 2.1. This helps the detector to focus on artifacts related to the synthetic generation process avoiding content and coding related biases.
@@
  • We show that incorporating proper content-based augmentation leads to better-calibrated results. This ensures that in-lab performance more closely aligns with the expected performance on real-world images shared across social networks.
@@
  • We study the effect of different distribution shifts and show that by leveraging a pre-trained large model fine-tuned end-to-end on our dataset, we achieve a SoTA accuracy superior to 90% even on unseen generators.
@@
Figure 3: Overview of existing ($a$, $b$, $c$) and proposed ($d$) strategies for building an aligned training dataset. Some methods try to match synthetic images to the corresponding real images by using class-based generation ($a$) or text-to-image generation with real images’ descriptions ($b$). In ($c$) real images are fed to an autoencoder to generate a reconstructed fake with the same content. Differently from ($c$), in our approach a self-conditioned fake is generated using diffusion steps ($d$), and we also add a content augmentation step.
@@ 2 Related Work
Alternatively, some methods rely on datasets comprising images from a single diffusion-based generator, such as Latent Diffusion [10, 11, 7], Guided Diffusion [47] or Stable Diffusion [40, 23]. Prior work [11, 7] highlights the importance of aligning both training and test data in terms of semantic content. This choice allowed to better exploit the potential of fixed-pretraining CLIP features by strongly reducing the number of images needed for fine-tuning [11]. In addition, it has the key merit of reducing the dataset content bias, thus allowing for better quality training, and is also adopted in other approaches both during training [1, 3] and at test time to carry out a fairer evaluation [2].
@@
Table 1: This table provides an overview of the datasets used in our evaluation, including the number of real and fake images, the sources of the real data, and the number of generative models used to create the synthetic images.
@@ 3.1 Datasets
3.2 Metrics
Most work on GenAI image detection measure performance by means of threshold-independent metrics, such as Area Under the Curve (AUC) or average precision (AP). These metrics indicate ideal classification performance, however the optimal separating threshold is not known and, quite often, the balanced accuracy at a fixed threshold (e.g. $0.5$) remains low, especially when there are significant differences between training and testing distributions [44]. Some papers address this problem by adjusting the threshold through a calibration procedure, assuming access to a few images from the synthetic generator under evaluation [46, 10, 33]. In a realistic situation the availability of such calibration images cannot be guaranteed.
In this work, to provide a comprehensive assessment of performance, we use both AUC and Accuracy at $0.5$, in addition we compute the Expected Calibration Error (ECE) and the Negative Log-Likelihood (NLL). ECE measures the ability of a model to provide prediction probabilities well aligned with the true probabilities. More precisely, we use the Binary ECE, which is the weighted average of the differences between the actual probability and the predicted probability across different bins [35]. Then, we use the balanced Negative Log-Likelihood [36], which evaluates the similarity between the distribution of the model’s predictions and the actual data distribution, penalizing both low confidence in the correct class and overconfidence in incorrect ones. More details on these metrics can be found in the supplementary material.
Synthbuster New Generators WildRF AVG
@@ 3.2 Metrics
AUC / bAcc paired by text -- 96.9 / 56.9 99.5 / 78.1 78.7 / 50.1 98.8 / 56.6 91.3 / 51.1 94.6 / 51.5 96.6 / 66.9 97.8 / 72.5 84.4 / 67.2 96.1 / 68.1 93.5 / 61.9
reconstructed -- 100. / 99.8 100. / 100. 81.1 / 52.2 99.1 / 75.8 96.7 / 62.1 98.2 / 64.0 99.7 / 88.8 98.9 / 97.5 77.8 / 75.5 94.8 / 91.1 94.6 / 80.7
reconstructed inpainted 100. / 99.9 100. / 99.9 90.1 / 62.0 99.2 / 79.5 98.1 / 77.2 95.8 / 69.1 99.4 / 89.1 98.8 / 95.9 78.7 / 76.1 94.9 / 90.3 95.5 / 83.9
self-conditioned -- 99.9 / 97.2 100. / 99.2 90.4 / 58.1 98.9 / 76.1 99.4 / 89.7 95.4 / 59.4 100. / 98.4 95.4 / 86.6 75.7 / 66.3 91.4 / 82.7 94.7 / 81.4
@@
" inpainted 100. / 99.4 100. / 99.6 96.7 / 77.8 99.4 / 92.8 99.9 / 99.2 98.7 / 87.5 99.9 / 98.4 98.9 / 94.4 89.4 / 81.2 97.5 / 92.0 98.0 / 92.2
" inpainted+ 99.9 / 98.8 100. / 99.7 99.7 / 95.9 99.6 / 96.8 100. / 99.6 97.9 / 85.3 99.3 / 95.1 98.0 / 95.0 96.0 / 89.8 99.4 / 96.5 99.0 / 95.2
" inpainted++ 100. / 99.6 100. / 99.8 99.7 / 95.7 99.9 / 98.2 100. / 99.7 99.3 / 92.3 99.9 / 98.9 99.0 / 95.6 95.8 / 86.3 99.7 / 97.4 99.3 / 96.4
@@
NLL / ECE paired by text -- 1.96 / .418 0.72 / .218 3.60 / .496 1.71 / .416 2.95 / .484 2.62 / .473 1.42 / .327 1.00 / .273 1.77 / .324 1.31 / .310 1.91 / .374
reconstructed -- 0.00 / .003 0.00 / .001 3.33 / .469 0.83 / .240 1.56 / .368 1.46 / .354 0.38 / .115 0.13 / .025 1.20 / .192 0.43 / .082 0.93 / .185
reconstructed inpainted 0.01 / .008 0.01 / .008 1.16 / .353 0.40 / .197 0.51 / .222 0.79 / .290 0.23 / .107 0.12 / .032 0.73 / .153 0.29 / .066 0.42 / .144
self-conditioned -- 0.08 / .031 0.02 / .008 1.48 / .399 0.59 / .234 0.27 / .114 1.22 / .379 0.04 / .021 0.31 / .084 0.92 / .237 0.40 / .078 0.53 / .158
@@
" inpainted 0.02 / .016 0.02 / .017 0.49 / .199 0.18 / .085 0.04 / .025 0.27 / .117 0.05 / .021 0.16 / .070 0.39 / .059 0.19 / .033 0.18 / .064
" inpainted+ 0.04 / .008 0.01 / .006 0.12 / .044 0.10 / .037 0.02 / .011 0.40 / .149 0.14 / .044 0.17 / .032 0.25 / .043 0.11 / .028 0.14 / .040
" inpainted++ 0.02 / .013 0.01 / .011 0.10 / .045 0.06 / .031 0.02 / .019 0.19 / .084 0.04 / .014 0.14 / .039 0.28 / .075 0.09 / .047 0.10 / .038
Table 2: Ablation study. We compare several forms of content alignment and content augmentation. Performance are in terms of AUC/Accuracy (top) and ECE/NLL (bottom). Note that all variants share a standard augmentation (blurring + JPEG compression) as proposed in [46]. For content alignment we consider image pairing strategies described in Fig. 3: text-driven generation, reconstruction through autoencoder, and our proposal using self-conditioned images. For the last solution we test several forms of augmentation: a standard cutmix/mixup, and three proposed strategies based on inpainting described in Sec. 4.2. Bold underlines the best performance for each column with a margin of 1%.
4 Proposed Method
To realize and test our bias-free training paradigm we:
  • build a dataset consisting of real and generated fake images, where the latter are well aligned with their real counterparts but include the forensic artifacts of the diffusion-based generation process. The dataset is created starting from the images collected from the training set of MS-COCO dataset [28], for a total of 51,517 real images. It is then enriched through several forms of augmentation, including locally inpainted images, and comprises eventually 309,102 generated images.
  • use this aligned dataset to fine-tune end-to-end a Vision Transformer (ViT)-based model. Specifically, we adopt a variant of the ViT network proposed in [13] with four registers and use the pretraining based on the self-supervised learning method DINOv2 [34]. During training, we avoid resizing the image and rely on large crops of $504\times 504$ pixels. At inference time, we also extract crops of the same dimension from the image (if the image is larger we average the results of multiple crops).